A 38.0 L gas tank at 35 c has nitrogen at a pressure of 4.65 atm. The contents of the tank are transferred without loss to an evacuated 55.0 L tank in a cold room where the temperature is 4 c. What is the pressure in the tank?
Answer #1

P1V1/T1 = P2V2/T2

4.65 x 38 /308 = P2 x 55/277

P2 = 4.65 x 38 x 277/(55 x 308) = 2.89 atm